The sum of 4 consecutive integers is 26. What is the product of the 4 numbers?
Answer by mathitutor(25)

Assume that the first number is x
x + x + 1 + x + 2 + x + 3 = 26
=> 4x + 6 = 26
=> 4x = 26 - 6
=> 4x = 20
=> x = 20/4
=> x = 5
Hence the numbers are: 5, 6, 7, 8
Product of these numbers = 5 x 6 x 7 x 8 = 1680